# TRAIL: Trace Reasoning and Agentic Issue Localization

### v1

@@ Abstract
The increasing adoption of agentic workflows across diverse domains brings a critical need to scalably and systematically evaluate the complex traces these systems generate. Current evaluation methods depend on manual, domain-specific human analysis of lengthy workflow traces—an approach that does not scale with the growing complexity and volume of agentic outputs. Error analysis in these settings is further complicated by the interplay of external tool outputs and language model reasoning, making it more challenging than traditional software debugging. In this work, we (1) articulate the need for robust and dynamic evaluation methods for agentic workflow traces, (2) introduce a formal taxonomy of error types encountered in agentic systems, and (3) present a set of 148 large human-annotated traces (TRAIL) constructed using this taxonomy and grounded in established agentic benchmarks. To ensure ecological validity, we curate traces from both single and multi-agent systems, focusing on real-world applications such as software engineering and open-world information retrieval. Our evaluations reveal that modern long context LLMs perform poorly at trace debugging, with the best gemini-2.5-pro model scoring a mere 11% on TRAIL. Our dataset and code are made publicly available to support and accelerate future research in scalable evaluation for agentic workflows [^1] https://huggingface.co/datasets/PatronusAI/TRAIL.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has catalyzed the development of complex agentic systems capable of automating difficult, multi-step tasks across various domains such as software engineering and multi-hop IR Ma et al. (2023); OpenAI (2024); Nguyen et al. (2024); Wang et al. (2025a). Unlike traditional generative models, agents can interact with diverse tools and dynamically navigate uncertain environments, often with minimal human supervision Wang et al. (2024a). This escalation of system complexity has leads to proposals for more challenging and multifaceted evaluation processes Nasim (2025) and has led to the adoption of LLMs as evaluators and critics of these agentic systems Zheng et al. (2023); Chen et al. (2024); Kim et al. (2024); Zhu et al. (2025); Deshpande et al. (2024a).
@@
However, as multi-agent systems become integral to real-world workflows, evaluating and debugging their performance remains a significant challenge. Agentic non-determinism Laban et al. (2025); Patronus AI (2025) and multi-step task solving Mialon et al. (2023); Yao et al. (2024) demand greater observability than the simple end-to-end evaluations offered by existing benchmarks Kapoor et al. (2024a); Zhuge et al. (2024); Moshkovich et al. (2025); Cemri et al. (2025). Such complex environments require granular taxonomies and well-annotated traces that can serve as references for debugging and root-cause analysis of agent behaviors Cemri et al. (2025). When creating taxonomies and benchmarks to test and improve agents, we must ensure these are grounded in real-world applications, following principles established in prior work Bowman and Dahl (2021); Liu et al. (2024b). Previous agent trace analysis frameworks have primarily focused on parsed traces containing unstructured text Cemri et al. (2025), which do not adequately represent common agent framework outputs that generate structured traces logged in standardized formats like opentelemetry OpenTelemetry (2025). We have observed that handling structured data remains challenging for LLMs Guo et al. (2023); Sui et al. (2024), an observation corroborated by previous research on automated software engineering trace analysis  Roy et al. (2024a); Ma et al. (2024b). These limitations highlight the need for new approaches specifically designed for structured agentic traces. To address these challenges and facilitate the analysis and evaluation of structured agentic executions, we propose a formal error taxonomy, shown in Figure 3, that promotes granular failure diagnosis. We also present a carefully curated, turn-level annotated trace dataset called TRAIL (Trace Reasoning and Agentic Issue Localization), which demonstrates the validity and practical utility of our proposed taxonomy.
@@
In our work, we utilize and build on SWE-Bench Jimenez et al. (2024); Aleithan et al. (2024) and GAIA Mialon et al. (2023) while addressing three major shortcomings inherent to previous automatic agent evaluation paradigms. Firstly, we aim to replace end to end analysis of agents Mialon et al. (2023); Wang et al. (2025a); Jimenez et al. (2024); Huang et al. (2024) with a benchmark containing step-level analysis of traced agentic workflows. Secondly, we address the need for grounding in real scenarios by producing opentelemetry-based structured traces that span beyond present model context length limits. Finally, as compared to benchmarks focused only on agentic reasoning and coordination Cemri et al. (2025); Kokel et al. (2025), TRAIL displays its validity through addition of finer, more aligned system execution failures and planning error categories such as API errors and Task Orchestration Errors to our taxonomy. Such categories are not only relevant to model developers but also to users and engineers optimizing single and multi-agent AI applications. The contributions of our work are as follows:
@@
    1. (a)
    2. (b)
      Solving TRAIL requires a significant fraction of the maximum input length of LLMs (or exceeds it), as well as requires generating significant fraction of their maximum output (See Table Table 2)
    3. (c)
@@ 3.1 Reasoning Errors
The generation of factually incorrect or nonsensical content is a pervasive issue with LLMs that extends to agents Huang et al. (2025); Ji et al. (2023). Text-only hallucinations manifest as deviations from factual reality or fabricated textual elements, such as ungrounded statements misaligned with established world knowledge Ji et al. (2023). On the other hand, tool-related hallucinations occur when agents fabricate tool outputs or misunderstand tool capabilities Zhang et al. (2024b). This can involve inventing results supposedly produced by a tool or claiming non-existent functionalities Xu et al. (2024)
@@
Recent years have observed rise in retrieval augmented generation, a process through which relevant data is retrieved based on similarity to a query and reasoned over Hu and Lu (2024); Gao et al. (2025). In their work Xu et al. (2025); Su et al. (2025) show that LLMs are poor at reasoning over retrieved instances for different tasks . The problems in information processing can be categorized into two subcategories: poor information retrieval and misinterpretation of reasoned outputs. Poor information retrieval Wu et al. (2024) is a critical problem since incorrect or irrelevant queries lead to redundancy in agent systems Wu et al. (2024) and can potentially lead to content overloading problems as observed with chain-of-thought reasoning models Stechly et al. (2024). Similarly, misinterpretation of retrieved context (Tool output MisinterpretationKarpinska et al. (2024); Wang et al. (2024b) may lead to incorrectness of a smaller, local task which propagates through the multiple steps of agent reasoning, leading to incorrectness or inefficiencies.
@@
At a step-level, task misunderstanding can stem from ambiguity in the input prompt, unclear instructions, or from the LLMs inability to differentiate the instructions in the data from the instructions in its prompt Zverev et al. (2024). Detecting if an agent has misunderstood a problem (Incorrect Problem ID) involves analyzing agent’s path trajectory Yuan et al. (2024) which, when combined with large contexts, makes error localization difficult. Hence, reliably detecting misinterpretation is essential for agent improvement.
Another important aspect of decision making in agent workflows involves selection of correct tools at every step Qin et al. (2023). Since plan-optimality leads to cost minimization and efficiency Yehudai et al. (2025), selection of the correct tool is essential. However, this selection is highly dependent on the task, available tools and contextual knowledge about the system. Hence, we include Tool Selection Error as a sub-category under Decision Making.
@@
LLMs have been shown to incorrectly format structured outputs Shorten et al. (2024); Liu et al. (2024a) and since most tool calls are made either via JSON responses or through code, correct formatting of outputs becomes necessary. To address this, we add the Formatting Errors subcategory to our taxonomy. On the other hand, works such as  White et al. (2024); Heo et al. (2024) show that LLMs are poor at instruction following when provided with complex or ambiguous instructions. To address such non-compliant outputs, we add the Instruction Non-compliance subcategory to our taxonomy.
@@ 3.2 System Execution Errors
Incorrect configuration of the agentic environment can lead to failures and restriction of agentic capabilities Hu et al. (2024a). A core example of an agentic configuration issue is Incorrect Tool Definition. In their recent work, Fu et al. (2024) show that agents can be tricked into using tools based on incorrect definitions or obfuscations in prompt which makes this subcategory a concern for both security and reliability of the system. Furthermore, poor configuration of environment variables (Environment Setup Errors) such as missing API keys, incorrect file access permissions can lead to unexpected failures in following correct reasoning paths generated by the model.
@@
Since agentic systems combine LLMs with software tools, issues with tool use or incorrect tool implementation can show up as errors during the workflow. Recently, agentic frameworks have seen increased adoption of remote tool access through protocols like Model Context Protocol Anthropic (2025) which highlights the need for capturing and categorizing API failures so that failures can be quickly reported to engineers building these remote tools Shen (2024). Additionally, Milev et al. (2025) also show that runtime errors with agentic tools are an underexplored problem. To capture these runtime errors with API tools, we add the most common errors (in addendum to Liu et al. (2023a)): Rate Limiting (such as 429 error), Authentication Errors (such as 401 and 403 errors), Service Errors (such as 500 errors) and Resource Not Found Errors (such as 404 error).
@@
Resource management is key for agents that have access to operating system tools like python interpreters or terminal access tools. In such scenarios, poor task understanding and planning in agents can surface vulnerabilities in systems. One prime example of such vulnerability is the exhaustion of resources allocated to the agent Ge et al. (2023) (Resource Exhaustion) or infinite loops Zhang et al. (2024a) (Timeout Issues) which can lead to cases of compute memory overflow or system overloads. Early detection of such errors is essential to prevent infrastructure collapse.
@@ 3.3 Planning and Coordination Errors
With increased adoption of planning and reasoning stages in agentic workflows Yao et al. (2023); Ke et al. (2025), reasoning over long contexts is an essential task for agents. In such scenarios, maintaining episodic and semantic Zhang et al. (2024c) context for information becomes necessary for an agent to improve. For the creation of this taxonomy, we label the context and instruction retention errors as Context Handling Failures. Tool call repetition Kokane et al. (2024) (Resource Abuse) is another clear failure of planning, context management and tool understanding in agents which we capture in our taxonomy.
@@
Environmental misconfigurations or LLM hallucinations can act as distractions in agentic systems. Poor recoverability from such distractions can lead to failures in task completion and goal deviation Ma et al. (2024a). This error is exacerbated with multi-agent systems and introduction of sub-tasks, making proper task orchestration an important aspect of agentic success. Hence, we include Goal Deviation and Task Orchestration Errors in our taxonomy.
GAIA SWE Bench
@@
Table 1: Performance on GAIA and SWE Bench. Models marked with * have reasoning set to "high"; † indicates 1M+ token context window. Insufficient context length is indicated by CLE. Pearson correlation between overall human and generated scores are represented in the $\rho$ column.
4 TRAIL Benchmark
While existing agent trace evaluation benchmarks focus on parsed traces and unstructured text Cemri et al. (2025), applications utilizing agentic frameworks  [^2] https://github.com/huggingface/smolagents, [^3] https://github.com/pydantic/pydantic-ai, [^4] https://github.com/langchain-ai/langchain tend to produce structured traces that are logged using standardized opentelemetry formats such as openinference Arize AI (2025); Moshkovich et al. (2025). TRAIL, following this standardization, is a benchmark aimed to evaluate LLM capabilities to analyze and evaluate these long, structured agentic executions. TRAIL follows our fine grained taxonomy and contains a total of 148 carefully annotated agent execution traces. TRAIL uses text-only data instances from the GAIA Mialon et al. (2023) and SWE Bench Lite Jimenez et al. (2024) datasets, spanning multiple information retrieval and software engineering bug fixing tasks. The TRAIL dataset contains a total of 841 annotated errors, averaging at 5.68 errors per trace.
4.1 Goals and Design Choices
Core Agent Task
We aim to showcase realistic agentic workflows and so we target two widely adopted agentic datasets, the GAIA benchmark Mialon et al. (2023), an open world search task, and the SWE-Bench-Lite Jimenez et al. (2024) dataset, for locating problems in Github repositories and creating fixes for them. We select these datasets due to their challenging nature, necessity for environment, search space exploration, and a good alignment with the taxonomy.
Agent Orchestration
Liu et al. (2023b) first presented a standardized hierarchical method of orchestrating agents, derivatives of which are actively adopted by several works Zhao et al. (2024, 2025). We closely follow this hierarchical structure and adopt the Hugging Face OpenDeepResearch agent Hugging Face (2024) for creating traces for the GAIA benchmark. We select the state-of-the-art o3-mini-2025-01-31 OpenAI (2025d) and assign it as the backbone model for the manager and search agents respectively because of its strong tool use and planning ability as showcased by Phan et al. (2025). For more information, refer to Appendix subsection A.6.
On the other hand, to explore single-agent planning errors and elicit context handling errors for the SWE-Bench split, we use a CodeAct agent Wang et al. (2024c) and provide it access to a sandboxed environment and a python interpreter and the gitingest [^5] https://github.com/cyclotruc/gitingest library. We select claude-3-7-sonnet-20250219 as the backbone model due to its strong performance on software engineering tasks Anthropic (2025). To further organically introduce errors into this agent pipeline, we add instructional constraints such as output text length limits and force exploration through prompts. The complete prompt can be found in the Appendix subsection A.7.
Workflow Tracing
To ensure compatibility of this dataset with real world tracing and observability software, all traces are collected with the help of opentelemetry OpenTelemetry (2025), specifically, its most widely adopted open-source derivative compatible with agents, the openinference standard  Arize AI (2025) following Moshkovich et al. (2025).
@@ 4.2 Data Annotation and Validation
We select four expert annotators with a background in software engineering and log debugging to annotate our agent traces. Furthermore, to ensure thoroughness and completeness of the annotation, we assign a separate set of 63 traces to calculate and analyze agreement. Since these traces are large and sometimes extend well beyond the maximum context length of most LLMs (as we detail in §5.1.1), we perform four independent rounds of verification with a set of four ML researchers to ensure high quality. During curation, annotators are asked to first iterate over each LLM and tool span and annotate each one individually and in context to the previous spans based on our taxonomy. For each span, the annotators mark the span ID, error category type, evidence, description and impact level (Low/Medium/High) associated with the error if it exists. At the end, the annotators are asked to rate the overall trace based on instruction adherence, plan optimality, security and reliability (definitions in the Appendix § A.3.1). On an average, a single trace of GAIA takes approximately 30 minutes to annotate whereas SWE Bench takes close to 40 minutes each. The additional verification step takes approximately 20 minutes for each trace, making the overall annotation time for each trace close to 110 minutes for GAIA and 120 minutes for SWE-Bench-Lite [^6] We did not explore and verify information (web-based or otherwise) from contents external to the trace because our baseline models are not expected to do so. Verifying such information will add more time to this estimate.. For the SWE Bench split, we reviewed 30 traces containing a total of 444 annotated spans, of which 25 spans (5.63%) were modified by reviewers. The top three error categories revised were Resource Abuse (33.33%), Language-only Hallucinations (20.83%) and Tool-related Hallucinations (12.5%). Alternately, we reviewed 33 traces with a total of 697 annotated spans and modified 37 spans (5.31%) after revision. The most commonly revised error types were Language-only Hallucinations (23.08%), Resource Abuse(19.23%) and Poor Information Retrieval(19.23%). This displays high level of interannotator agreement during data curation.

### v2

@@ Abstract
The increasing adoption of agentic workflows across diverse domains brings a critical need to scalably and systematically evaluate the complex traces these systems generate. Current evaluation methods depend on manual, domain-specific human analysis of lengthy workflow traces—an approach that does not scale with the growing complexity and volume of agentic outputs. Error analysis in these settings is further complicated by the interplay of external tool outputs and language model reasoning, making it more challenging than traditional software debugging. In this work, we (1) articulate the need for robust and dynamic evaluation methods for agentic workflow traces, (2) introduce a formal taxonomy of error types encountered in agentic systems, and (3) present a set of 148 large human-annotated traces (TRAIL) constructed using this taxonomy and grounded in established agentic benchmarks. To ensure ecological validity, we curate traces from both single and multi-agent systems, focusing on real-world applications such as software engineering and open-world information retrieval. Our evaluations reveal that modern long context LLMs perform poorly at trace debugging, with the best gemini-2.5-pro model scoring a mere 11% on TRAIL. Our dataset and code are made publicly available to support and accelerate future research in scalable evaluation for agentic workflows [^1] Hidden for double-blind review.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has catalyzed the development of agentic systems capable of automating difficult, multi-step tasks across various domains such as software engineering and multi-hop IR Ma et al. (2023); OpenAI (2024); Nguyen et al. (2024); Wang et al. (2025a). Unlike traditional generative models, agents can interact with diverse tools and dynamically navigate environments, often with minimal human supervision Wang et al. (2024a). This escalation of system complexity demands more challenging and multifaceted evaluation processes Nasim (2025) and has led to the adoption of LLMs as evaluators for such agentic systems Zheng et al. (2023); Chen et al. (2024); Kim et al. (2024); Zhu et al. (2025); Deshpande et al. (2024a).
@@
However, as multi-agent systems scale and become integral to real-world workflows, evaluating and debugging their performance remains a significant challenge. Agentic non-determinism Laban et al. (2025); Patronus AI (2025) and multi-step task solving Mialon et al. (2023); Yao et al. (2024) demand greater observability than the simple end-to-end evaluations offered by existing benchmarks Kapoor et al. (2024a); Zhuge et al. (2024); Moshkovich et al. (2025); Cemri et al. (2025). Such complex environments require granular taxonomies and well-annotated traces that can serve as references for debugging and root-cause analysis of agent behaviors Cemri et al. (2025). When creating taxonomies and benchmarks to test and improve agents, we must ensure these are grounded in real-world applications and are not centered around dummy data Bowman and Dahl (2021); Liu et al. (2024b). Previous agent trace analysis frameworks have primarily focused on parsed traces containing unstructured text Cemri et al. (2025), which do not adequately represent common agent framework outputs that generate structured traces logged in standardized formats like opentelemetry OpenTelemetry (2025). Additionally, as observed by  Guo et al. (2023); Sui et al. (2024), handling structured data remains challenging for LLMs, an observation corroborated by previous research on automated software engineering trace analysis  Roy et al. (2024a); Ma et al. (2024b). These limitations highlight the need for new approaches specifically designed for structured agentic traces. To address these challenges and facilitate the analysis and evaluation of agentic executions, we propose a formal error taxonomy, shown in Figure 3, that promotes granular failure diagnosis. We also present a carefully curated, turn-level annotated trace dataset called TRAIL (Trace Reasoning and Agentic Issue Localization), which demonstrates the validity and practical utility of our proposed taxonomy.
@@
In our work, we utilize and build on SWE-Bench Jimenez et al. (2024); Aleithan et al. (2024) and GAIA Mialon et al. (2023) while addressing three major shortcomings inherent to previous automatic agent evaluation paradigms. Firstly, we aim to replace end to end analysis of agents with a benchmark containing step-level analysis of traced agentic workflows. Secondly, we address the need for grounding in real scenarios by producing opentelemetry-based structured traces that span beyond present model context length limits. Finally, as compared to benchmarks focused only on agentic reasoning and coordination Cemri et al. (2025); Kokel et al. (2025), TRAIL focuses on validity through addition of finer, more aligned system execution failures and planning error categories such as API errors and Task Orchestration Errors to our taxonomy. Such categories are not only relevant to model developers but also to users and engineers optimizing single and multi-agent AI applications. The contributions of our work are as follows:
@@
      Solving TRAIL requires a significant fraction of the maximum input length of LLMs (or exceeds it), as well as requires generating significant fraction of their maximum output (See Table 2, Figure 5)
@@
Figure 2: TRAIL trace’s span structure and error examples
@@ 3.1 Reasoning Errors
LLMs frequently generate factually incorrect or nonsensical content, a problem that also affects agents Huang et al. (2025); Ji et al. (2023). Text-only hallucinations include fabricated or ungrounded statements that conflict with real-world knowledge Ji et al. (2023). In contrast, Tool-related hallucinations arise when agents invent tool outputs or misunderstand tool functions, such as fabricating results or claiming nonexistent capabilities Zhang et al. (2024b); Xu et al. (2024).
@@
Retrieval-augmented generation, which retrieves and reasons over data relevant to a query, has become increasingly popular Hu and Lu (2024); Gao et al. (2025). However, recent work Xu et al. (2025); Su et al. (2025) shows that LLMs and agents often struggle to reason effectively over retrieved information. These issues can be grouped into two main types: poor information retrieval and misinterpretation of outputs. Poor information retrieval Wu et al. (2024) can introduce redundancy and content overload Stechly et al. (2024), while misinterpretation of retrieved context (Tool output MisinterpretationKarpinska et al. (2024); Wang et al. (2024b) may cause errors that propagate throughout an agent’s reasoning process, leading to broader incorrectness or inefficiencies.
@@
Task misunderstanding at the step level often arises from ambiguous prompts, unclear instructions, or an LLM’s inability to distinguish between prompt and data instructions Zverev et al. (2024). Detecting such misunderstandings (Incorrect Problem ID) requires analyzing an agent’s path, which is challenging in large contexts Yuan et al. (2024) and reliable detection of these errors is crucial for agent improvement. Furthermore, effective decision making in agent workflows also depends on selecting the appropriate tool at each step Qin et al. (2023). Because optimal planning and tool selection reduces cost and increases efficiency Yehudai et al. (2025), we place Tool Selection Error under Decision Making.
@@
LLMs often produce incorrectly formatted structured outputs Shorten et al. (2024); Liu et al. (2024a), which is problematic for tool calls that need precise JSON or code formatting. To capture this, our taxonomy includes Formatting Errors. Similarly, LLMs frequently struggle following complex/ambiguous instructions White et al. (2024); Heo et al. (2024), hence we subcategorize Instruction Non-compliance.
@@ 3.2 System Execution Errors
Incorrect agentic environment configuration can cause failures and limit agent capability Hu et al. (2024a). One key issue is Incorrect Tool Definition, as shown by Fu et al. (2024), agents can be misled by inaccurate or obfuscated tool definitions in prompts, posing security and reliability risks. Additionally, poor setup of environment variables (Environment Setup Errors), e.g., missing API keys or incorrect file permissions, can cause unexpected failures and disrupt reasoning paths.
@@
As agentic systems combine LLMs with software tools, tool usage or implementation errors can disrupt workflows. With the rise of remote tool access via protocols like MCP Anthropic (2025), capturing and categorizing API failures is increasingly important for prompt reporting to tool developers Shen (2024). Runtime errors involving agentic tools remain underexplored Milev et al. (2025), so we specifically include the most common API tool errors in our taxonomy: Rate Limiting (429), Authentication Errors (401, 403), Service Errors (500), and Resource Not Found (404) Liu et al. (2023a).
@@
Resource management is crucial for agents using operating system tools like interpreters or terminals. Poor task planning can expose vulnerabilities, such as Resource Exhaustion from overallocation Ge et al. (2023) or Timeout Issues from infinite loops Zhang et al. (2024a), potentially causing memory overflows or system overloads. Early detection of these errors is vital to prevent infrastructure failures.
@@ 3.3 Planning and Coordination Errors
As planning and reasoning become integral to agentic workflows Yao et al. (2023); Ke et al. (2025), agents must manage long-term context, including episodic and semantic information Zhang et al. (2024c). In our taxonomy, we categorize failures in context or instruction retention as Context Handling Failures. Additionally, repeated tool calls Kokane et al. (2024) (Resource Abuse) reflect shortcomings in planning, context management, and tool use, which our taxonomy also captures.
@@
Environmental misconfigurations or LLM hallucinations can distract agentic systems, and poor recovery from such distractions often leads to goal deviation Ma et al. (2024a). These issues are amplified in multi-agent setups with sub-tasks, making effective task orchestration crucial. Therefore, we include Goal Deviation and Task Orchestration Errors in our taxonomy.
(a) Error Category Annotations in TRAIL
(b) Distribution of Errors per Trace
Figure 3: TRAIL Dataset Statistics
4 TRAIL Benchmark
TRAIL is a benchmark aimed to evaluate LLM capabilities to analyze and evaluate long, structured, opentelemetry standardized agentic executions. TRAIL follows our fine grained taxonomy and contains 148 carefully annotated agentic traces. The dataset uses text-only data instances from the GAIA Mialon et al. (2023) and SWE Bench Lite Jimenez et al. (2024) datasets, spanning multiple information retrieval and software bug fixing tasks. It contains a total of 841 annotated errors, averaging at 5.68 errors per trace Figure 3.
4.1 Goals and Design Choices
Core Agent Task
We aim to showcase realistic agentic workflows and so we target two widely adopted agentic datasets, the GAIA benchmark Mialon et al. (2023), an open world search task, and the SWE-Bench-Lite Jimenez et al. (2024) dataset, for locating and fixing issues in Github repositories. We select these datasets due to their challenging nature and necessity for environment and search space exploration.
Agent Orchestration
Liu et al. (2023b) first presented a standardized hierarchical method of orchestrating agents, derivatives of which are actively adopted by several works Zhao et al. (2024, 2025). We closely follow this hierarchical structure and adopt the Hugging Face OpenDeepResearch agent Hugging Face (2024) for creating traces for the GAIA benchmark. We select the state-of-the-art o3-mini-2025-01-31 OpenAI (2025d) and assign it as the backbone model for the manager and search agents respectively because of its strong tool use and planning ability as showcased by Phan et al. (2025). For more information, refer to §A.10.
Parallelly, to explore single-agent planning errors and elicit context handling errors for the SWE-Bench split, we use a CodeAct agent Wang et al. (2024c) and provide it access to a sandboxed environment, a python interpreter and the gitingest [^2] https://github.com/cyclotruc/gitingest library. We select claude-3-7-sonnet-20250219 as the backbone model due to its strong performance on software engineering tasks Anthropic (2025). To further organically introduce errors into this agent system, we add instructional constraints such as output length limits and force exploration via prompts. The complete prompt is at §A.12.
Workflow Tracing
To ensure compatibility of this dataset with real world tracing and observability software, all traces are collected via opentelemetry OpenTelemetry (2025), specifically, its most widely adopted open-source derivative compatible with agents, the openinference standard  Arize AI (2025) as adopted by Moshkovich et al. (2025).
TRAIL (GAIA) TRAIL (SWE Bench)
@@ 4.1 Goals and Design Choices
Table 1: Performance across LLMs for Error Categorization & Localization on TRAIL (GAIA) and TRAIL (SWE Bench). Models marked with * have reasoning set to "high"; † indicates 1M+ token context window. Insufficient context length is marked as CLE. Pearson correlation b/w overall human and generated scores is shown under $\rho$ [^4] All reported results are an average of three runs.
@@ 4.2 Data Annotation and Validation
We selected four annotators with expertise in software engineering and log debugging to label our agent traces. To assess agreement, a separate set of 63 traces was assigned. Results based on these indicate high inter-annotator agreement during curation. We defer details of our complete annotation and agreement measuring processes and actual numbers from them to §A.7.